OpenStudy (anonymous):

Does the following set of ordered pairs satisfy a linear function?(1,-5),(2,-3),(3,-1),(4,1),(5,3)

OpenStudy (science0229):

If all of the points are on a linear function, their slopes should be all same. If there are 2 points (a,b) and (c,d), the slope is \[\frac{ d-b }{ c-a }\]. So for the first 2 points, the slope is \[\frac{ -3--5 }{ 2-1 }=2\] Do the same for the rest to see if their slope is 2. If one of the pairs doesn't make the slope of 2, the points are not on a linear function.

OpenStudy (science0229):

Skipping all the calculations, all of the points make the slope of 2, so all of them are on a linear function.